OpenStudy (anonymous):

use implicit differentation given x^3-2xy+y^2=4 points (1,-1)

OpenStudy (bahrom7893):

3*x^2 - 2(x*y' + y) + 2y*y' = 0

OpenStudy (bahrom7893):

At (1,-1) 3 * 1^2 - 2(1 * y' + (-1)) + 2 (-1) * y' = 0

OpenStudy (bahrom7893):

3 - 2(y' -1) -2y' = 0

OpenStudy (bahrom7893):

3 - 2y' + 2 - 2y' = 0 -4y' = -5 y' = 5/4

OpenStudy (bahrom7893):

hope this helped lol

OpenStudy (anonymous):

yes :=)